OpenStudy (anonymous):

evaluate the integral dx/sq root of x(1+x) dx.. can u detail it for me please

OpenStudy (anonymous):

\[\sqrt{x}(x+1)\] ? 1. Distribute x^(1/2) get (x^(1/2)+x^(3/2)) 2. Integrate normally get (2/3)x^(3/2)+(2/5)x^(5/2) 3. Check via derive get X^(3/2)+sqrt(x) = sqrt(x)(X+1) check

OpenStudy (anonymous):

they gave me a different answer

OpenStudy (anonymous):

Hmm... possible lemme check it on my TI really quick I did this in my head.

OpenStudy (anonymous):

they said 2tan^-1sqroot x+c

OpenStudy (anonymous):

Ti is giving me .4x^(5/2)+.666667x^(3/2) which is approx what I answered. Are you sure you asked the right question or are looking at the right answer